# From Discrete to Continuous Binary Best-Response Dynamics: Discrete Fluctuations Almost Surely Vanish with Population Size

### v3

@@ Abstract
In binary decision-making, individuals often choose either the rare or the common action. In the framework of evolutionary game theory, the best-response update rule can be used to model this dichotomy. Those who prefer the common action are called coordinators, and those who prefer the rare one are called anticoordinators. A finite mixed population of the two types may undergo perpetual fluctuations, the characterization of which appears to be challenging. It is particularly unknown whether the fluctuations persist as population size grows. To fill this gap, we approximate the discrete finite population dynamics of coordinators and anticoordinators with the associated mean dynamics in the form of differential inclusions. We show that the family of the state sequences of the discrete dynamics for increasing population sizes forms a generalized stochastic approximation process for the differential inclusion. On the other hand, we show that the differential inclusions always converge to an equilibrium. This implies that the reported perpetual fluctuations in the finite discrete dynamics of coordinators and anticoordinators almost surely vanish with population size. The results encourage to first analyze the often simpler continuous-time mean dynamics of the discrete population dynamics as the continuous-time dynamics partly reveal the asymptotic behavior of the discrete dynamics.
@@ I Introduction
Humans face a variety of repetitive decision-making problems, such as whether to follow the fashion trends, get a flu shot, or sign a petition [1, 2]. In such two-option decision-making problems, individuals are often either conformists, also known as coordinators, those who adopt a decision whenever it has already been adopted by a particular population proportion of decision-makers, or nonconformists, also known as anticoordinators, those who go for a decision whenever it has been rejected by a particular population proportion [3, 4].
@@
In the framework of evolutionary game theory, the so-called best-response update rule captures the decision-making processes of coordinators and anticoordinators by assuming linear payoff functions. An increasing payoff function with respect to the commonality of the decisions makes the best-response update rule equivalent to the linear threshold model, where the individuals are coordinators with possibly unique thresholds [5]. A decreasing payoff function, on the other hand, models the anticoordinators.
@@
Analyzing these decision-making processes requires accounting for each individual’s actions, and consequently, a large body of research studied the steady state behavior of finite discrete populations of best-responders. The problem has been investigated in in various settings, ranging from well-mixed [6] to structured populations [7, 8, 4], and from time-invariant to time-varying thresholds [9]. Semi-tensor products were also utilized to analyze networked evolutionary games [10, 11] and logical dynamic games [12, 13] .
@@
A finite population of heterogeneous coordinators, who differ in perceiving the commonality, admits equilibria [14] and equilibrates under asynchronous activation sequence [15]. A finite population of heterogeneous anticoordinators either equilibrates or converges to a two-state cycle [16, 17]. As for finite populations of coordinating and anticoordinating agents, it was shown that the dynamics may undergo perpetual fluctuations [18, 17, 19]. The exact form and the conditions for the existence of the fluctuations appear to be an open problem.
@@
Researchers get around the challenges arising from the analysis of finite population dynamics by exploiting the associated deterministic mean dynamics [20, 21], where the population is assumed to be infinitely large. In this regard, most studies investigated the replicator dynamics which are the mean dynamics of the other popular update rule, imitation [22, 23, 24, 25]. Fewer studies considered the best-response mean dynamics, which are differential inclusions [26, 27, 28, 29, 30]. Some studies approximated the evolution of decisions in structured populations with ordinary differential equations [31, 32, 33].
@@
Although analysis of the mean dynamics is generally more straightforward compared to that of finite populations, the potential discrepancy between these two may question the validity of such an approximation. Hence, much effort has been devoted to connecting the behavior of finite populations with the associated mean dynamics as the population size grows [34, 35, 36, 37]. The results on infinite horizon behavior connect the Birkhoff center of the mean dynamics with the support of the stationary measures of the Markov chain corresponding to the finite population’s dynamics [35, 38].
@@
How can these results, which link the mean dynamics to finite population dynamics, be applied to the heterogeneous populations of coordinators and anticoordinators? To the best of our knowledge, no study investigated the asymptotic behavior of infinite heterogeneous populations of exclusively coordinators, exclusively anticoordinators, or a mixture of both. Even if there were such results, could they shed light on the discrete population dynamics?
@@
We provide the answer in this paper, namely by using the deterministic differential inclusions as an approximation tool to reveal the asymptotic behavior of the discrete population dynamics as the population size approaches infinity. First, through an intuitive example we show that the perpetual fluctuations do not scale as the population size grows. Second, we write the population dynamics as a Markov chain and obtain the associated continuous-time population dynamics–an upper semicontinuous differential inclusion. Third, we show that the collection of population dynamics Markov chains define generalized stochastic approximation processes for the continuous-time population dynamics [38]. This allows the use of the results in [38] to approximate the asymptotic behavior of the discrete population dynamics by finding the Birkhoff center of the continuous-time population dynamics, which we do in the last part of the paper.
@@
Our contribution is fourfold: i) We show that in the continuous-time population dynamics, the exclusively heterogeneous population of coordinators may admit two types of equilibrium points, clean-cut and ruffled coordinator-driven, where clean-cut equilibria are asymptotically stable and ruffled are unstable–Lemma 2 and Theorem 2. On the contrary, the exclusively heterogeneous population of anticoordinators admit only one globally asymptotically stable equilibrium point that could be either clean-cut or ruffled anti-coordinator driven–Lemma 2 and 1. These results are in line with the existing results on the steady-state behavior of finite heterogeneous populations of exclusively coordinators [15] and exclusively anticoordinators [16]. ii) We show that in the continuous-time population dynamics, the mixed heterogeneous population of coordinators and anticoordinators may admit all three types of equilibria, clean-cut, ruffled anticoordinator-driven, and ruffled coordinator-driven, where clean-cut and ruffled anticoordinator-driven equilibria are asymptotically stable, the other type is unstable, and the dynamics always converge to one of the equilibria–Lemma 2 and Theorem 2. iii) Consequently, we show that the reported perpetual fluctuations in the finite discrete dynamics of coordinators and anticoordinators [19] almost surely vanish with the population size –Theorem 3 and 4, which are the main result of this paper. iv) Out of the analysis of the continuous-time population dynamics, insightful hints on the asymptotic behavior of the finite populations were obtained, such as having the same equilibria and similar long-term behavior. Hence, in general for other population dynamics, it may prove useful to study the behavior of the continuous-time population dynamics, which is usually more straightforward, prior to the precise analysis of the finite population dynamics.
@@
The set of real, nonnegative real, integer, nonnegative integer, and positive integer numbers are respectively shown by $\mathbb{R}$, $\mathbb{R}_{\geq 0}$, $\mathbb{Z}$, $\mathbb{Z}_{\geq 0}$, and $\mathbb{N}$. With an abuse of notation, for positive integers $n$ and $m$, define $\frac{1}{m}\mathbb{Z}^{n}=\{\frac{1}{m}\bm{y}\mid\bm{y}\in\mathbb{Z}^{n}\}$. For each $i\in\mathbb{N}$, $[i]$ denotes $\{1,2,\ldots,i\}$. Boldface letters refer to vectors. The notation $\bm{1}$ refers to a vector of appropriate size with all elements equal to $1$. The $i^{\text{th}}$ element of the vector $\mathbf{q}$ is indicated by ${q}_{i}$. The calligraphic font $\mathcal{X}$ represents a set. A sequence of variables $x_{0},x_{1},x_{2},\ldots$ is represented by $\langle x_{k}\rangle{{{}_{k=0}^{\infty}}}$. The Euclidean norm of a vector $\bm{x}$ is denoted by $|\bm{x}|$. By the notations $[a_{1},b_{1}]-[a_{2},b_{2}]$ and $[a_{1},b_{1}]-{c}$ we mean $[a_{1}-a_{2},b_{1}-b_{2}]$ and $[a_{1}-c,b_{1}-c]$, respectively. The notation $\text{cl}(\mathcal{X})$ denotes the closure of the set $\mathcal{X}$. A set-valued map ${\bm{\mathcal{V}}}(\bm{x})$ from $\mathbb{R}^{n}$ to $\mathbb{R}^{n}$ is denoted by notation ${\bm{\mathcal{V}}}:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$.
@@ II Problem Formulation
We consider a well-mixed population of $\mathsf{N}$ agents, labeled by $1,2,\ldots,\mathsf{N}$, choosing repeatedly between two strategies $\mathtt{A}$ and $\mathtt{B}$ over a discrete time sequence $t\in\frac{1}{\mathsf{N}}\mathbb{Z}_{\geq 0}$. The time sequence is indexed by $k$ where $k=\mathsf{N}t$ (i.e., the index of time $\frac{k}{\mathsf{N}}$ is $k$). Each agent $i\in[\mathsf{N}]$ has a threshold $\tau(i)\in(0,1)$ and is either a coordinator (conformist) or an anticoordinator (a nonconformist). A coordinator $i$ (resp. an anticoordinator $i$) tends to choose strategy $\mathtt{A}$ whenever the population proportion of strategy-$\mathtt{A}$ players, excluding herself, denoted $x^{\mathsf{N}}_{{-i}}\in[0,1]$, does not fall short of (resp. does not exceed) her threshold, i.e., $x^{\mathsf{N}}_{{-i}}\geq\tau(i)$ (resp. $\leq\tau(i)$).
@@
Coordinators (resp. anticoordinators) who have the same threshold build up a subpopulation, and there all altogether $\mathsf{p}^{\prime}$ (resp. $\mathsf{p}$) subpopulations of coordinators (resp. anticoordinators). Coordinating (resp. anticoordinating) subpopulations are labeled in the ascending (resp. descending) order of their thresholds by $1,\ldots,\mathsf{p}^{\prime}$ (resp. $\mathsf{p}$), that is, $\tau^{\prime}_{1}<\tau^{\prime}_{2}<\ldots<\tau^{\prime}_{\mathsf{p}^{\prime}}$ (resp. $\tau_{1}>\tau_{2}>\ldots>\tau_{\mathsf{p}}$), where $\tau^{\prime}_{p}$ (resp. $\tau_{p}$) is the threshold of the $p^{\text{th}}$ coordinating (resp. anticoordinating) subpopulation. So the heterogeneity of the population is captured by the distribution of population proportions over the total $\mathsf{p}+\mathsf{p}^{\prime}$ subpopulations as $\bm{\rho}=({\rho}_{1},\ldots,{\rho}_{\mathsf{p}},{\rho}^{\prime}_{\mathsf{p}^{\prime}},\ldots,{\rho}^{\prime}_{1})^{\top}$ where ${\rho}^{\prime}_{j}$ (resp. ${\rho}_{j}$) represents the ratio of the number of individuals in coordinating (resp. anticoordinating) subpopulation $j$ to the population size $\mathsf{N}$. For convenience of vector indexing, for $j\geq\mathsf{p}+1$, we define ${\rho}_{j}={\rho}^{\prime}_{\mathsf{p}+\mathsf{p}^{\prime}+1-j}$. Note that there is no subpopulation of size zero, meaning that $\min_{j}\{\rho_{j}\}\geq\frac{1}{\mathsf{N}}$.
@@
According to the best-response update rule, the active agent chooses the strategy which maximizes her current payoff, i.e, $s_{i}(k+1)=\mathtt{A}$ if $u_{i}^{\mathtt{A}}(k)\geq u_{i}^{\mathtt{B}}(k)$ and $\mathtt{s}_{i}(k+1)=\mathtt{B}$, otherwise where $u_{i}^{\mathtt{j}}(k)$ is the payoff of agent $i$ for choosing strategy $\mathtt{j}\in\{\mathtt{A},\mathtt{B}\}$ at time index $k$. If the agents’ payoffs are a linear function of the population proportion of strategy-$\mathtt{A}$ players, update rules (1) and (2) are equivalent to the best-response update rule.
@@
where $\mathit{x}^{\prime\mathsf{N}}_{p}$ (resp. $\mathit{x}^{\mathsf{N}}_{p}$) represents the proportion of $\mathtt{A}$-players who belong to coordinating (resp. anticoordinating) subpopulation $p$. We again use the notation $x^{\mathsf{N}}_{p}=x^{\prime\mathsf{N}}_{\mathsf{p}+\mathsf{p}^{\prime}+1-p}$ for $p>\mathsf{p}$. By defining
@@
The evolution of the state defines the population dynamics and are governed by update rules (1), (2), and the activation sequence of the agents. More specifically, the activation sequence is generated by the sequence of random variables $\langle{A}_{k}\rangle{{{}_{k=0}^{\infty}}}$ where ${A}_{k}$ is the agent active at time index $k$ and takes values in $[\mathsf{N}]$ with the distribution $\mathbb{P}[{A}_{k}=i]=\frac{1}{\mathsf{N}}$. A realization of an activation sequence $\langle{A}_{k}\rangle_{k=0}^{\infty}$ is persistent [39], if for every agent $i$ and every $K>0$ there exists a time index $T>K$ such that $A_{T}=i$.
@@
The discrete best-response population dynamics are defined by the following discrete time stochastic equation for $k=0,1,\ldots$
@@
where ${P}_{k}$ and ${S}_{k}$ are scalar random variables with distributions $\mathbb{P}[{P}_{k}=p]=\rho_{p}$, $\mathbb{P}[{S}_{k}=1|{P}_{k}=p]=x^{\mathsf{N}}_{p}/\rho_{p}$, $\mathbb{P}[{S}_{k}=2|{P}_{k}=p]=1-x^{\mathsf{N}}_{p}/\rho_{p}$, $\mathbf{e}_{{P}_{k}}$ is ${P}_{k}$th standard basis vector in $\mathbb{R}^{\mathsf{p}+\mathsf{p}^{\prime}}$, and ${s^{*}}(p,\bm{x}^{\mathsf{N}},{s})$ is a function where ${s^{*}}(p,\bm{x}^{\mathsf{N}},{1})$ returns $1$ (resp. $2$) if $\mathtt{A}$ (resp. $\mathtt{B}$) is the preferred strategy of an $\mathtt{A}$-playing agent and ${s^{*}}(p,\bm{x}^{\mathsf{N}},{2})$ returns $1$ (resp. $2$) if $\mathtt{A}$ (resp. $\mathtt{B}$) is the preferred strategy of a $\mathtt{B}$-playing agent in subpopulation $p$ at population state $\bm{x}^{\mathsf{N}}$ :
@@
$=$ $\begin{cases}{1}&\text{if }(x^{\mathsf{N}}\leq{\tau_{p}}\text{ and }p\leq\mathsf{p}\text{ {and} }{S_{k}=2})\\ &\text{ or }(x^{\mathsf{N}}\geq\tau_{\mathsf{p}+\mathsf{p}^{\prime}+1-p}^{\prime}\text{ and }p>\mathsf{p}\text{ {and} }{S_{k}=2})\\ &\text{ or }(x^{\mathsf{N}}\leq{\tau_{p}}{+\frac{1}{\mathsf{N}}}\text{ and }p\leq\mathsf{p}\text{ {and} }{S_{k}=1})\\ &\text{ or }(x^{\mathsf{N}}\geq\tau_{\mathsf{p}+\mathsf{p}^{\prime}+1-p}^{\prime}{+\frac{1}{\mathsf{N}}}\text{ and }p>\mathsf{p}\text{ {and} }{S_{k}=1}),\\ {2}&\text{if }(x^{\mathsf{N}}>{\tau_{p}}\ \text{ and }p\leq\mathsf{p}\text{ {and} }{S_{k}=2})\\ &\text{ or }(x^{\mathsf{N}}<\tau_{\mathsf{p}+\mathsf{p}^{\prime}+1-p}^{\prime}\text{ and }p>\mathsf{p}\text{ {and} }{S_{k}=2})\\ &\text{ or }(x^{\mathsf{N}}>{\tau_{p}}{+\frac{1}{\mathsf{N}}}\ \text{ and }p\leq\mathsf{p}\text{ {and} }{S_{k}=1})\\ &\text{ or }(x^{\mathsf{N}}<\tau_{\mathsf{p}+\mathsf{p}^{\prime}+1-p}^{\prime}{+\frac{1}{\mathsf{N}}}\text{ and }p>\mathsf{p}\text{ {and} }{S_{k}=1}),\end{cases}$
@@
In Definition 1, the random variable ${P}_{k}$ is the subpopulation of the agent active at time index $k$, and the random variable ${S}_{k}$ equals $1$ (resp. $2$) if the strategy of the active agent at index $k$ equals $\mathtt{A}$ (resp. $\mathtt{B}$). Persistent realizations of the sequences of random variables $\langle P_{k}\rangle_{k=0}^{\infty}$ and $\langle S_{k}\rangle_{k=0}^{\infty}$ refer to their realizations under a persistent activation sequence. Function $s^{*}(p,\bm{x}^{\mathsf{N}},\cdot)$ returns $1$ if $\mathtt{A}$ is the preferred strategy of an agent in the subpopulation $p$ when the population is at state $\bm{x}^{\mathsf{N}}$, and it returns $2$ otherwise.
@@
In the context of public health, when a newly developed vaccine is introduced, some individuals choose to get vaccinated (i.e., strategy $\mathtt{A}$) as long as the vaccination coverage is below a certain threshold [40]. For them, this signals that not enough individuals are vaccinated, and they want to be immunized. Conversely, some individuals go for vaccination when a specific proportion of the population has already been vaccinated [41] as then they perceive the vaccine as safe or a societal norm. The first group can be thought of as anticoordinators, while the latter are coordinators. The perception of what proportion of the population needs to be vaccinated to be considered rare or common varies among individuals, resulting in a heterogeneous population.
What is the asymptotic behavior of the population dynamics as the population size approaches infinity? Does the population state reach an equilibrium point where no agent tends to switch her adopted strategy? Or do the population dynamics undergo perpetual fluctuations?
We define a fluctuation set as a non-empty subset of the state space that is (i) positively invariant, meaning that once the state enters the set it never leaves it under any activation sequence, (ii) it is minimal, meaning that it does not admit a positively invariant proper subset, implying that once the state enters it, every state in the set will be visited under a persistent activation sequence, and (iii) it is non-singleton, implying that it is not an equilibrium. In what follows, we provide the definition in terms of the dynamics in Definition 1.
@@
The dynamics of a finite mixed population consisting of both coordinating and anticoordinating subpopulations may undergo perpetual fluctuations [19], the characterization of which appears to be challenging and remain unsolved. Could the asymptotic behavior of discrete populations be revealed by investigating the associated mean dynamics, as the population size approaches infinity?

### v4

@@ Abstract
In binary decision-making, individuals often go for a common or a rare action. In the framework of evolutionary game theory, the best-response update rule can be used to model this dichotomy. Those who prefer the common action are called coordinators or conformists, and those who prefer the rare one are called anticoordinators or nonconformists. A finite mixed population of the two types may undergo perpetual fluctuations, the characterization of which appears to be challenging. It is particularly unknown whether the fluctuations persist as the population size grows. To fill this gap, we approximate the discrete population dynamics of coordinators and anticoordinators with the associated mean dynamics in the form of differential inclusions. We show that the family of state sequences of the discrete dynamics with increasing population sizes forms a generalized stochastic approximation process for the differential inclusion. On the other hand, we show that the differential inclusion always converges to an equilibrium. This implies that the reported perpetual fluctuations in the discrete dynamics of coordinators and anticoordinators almost surely vanish with population size. The results motivate analyzing the often simpler mean dynamics, which partly reveal the asymptotic behavior of the discrete dynamics.
@@ I Introduction
Humans face a variety of repetitive decision-making problems, such as whether to follow the fashion trends, get a flu shot, or sign a petition [cao2013fashion, blanton2001introduction]. In such two-option decision-making problems, individuals are often either coordinators, also known as conformists, those who adopt a decision whenever it has already been adopted by a particular population proportion of decision-makers, or anticoordinators, also known as nonconformists, those who go for a decision whenever it has been rejected by a particular population proportion [coordinationandanticoordination, arditt2].
@@
In the framework of evolutionary game theory, the so-called best-response update rule can capture the decision-making processes of coordinators and anticoordinators by assuming affine payoff functions. An increasing payoff function with respect to the commonality of the decisions makes the best-response update rule equivalent to the linear threshold model, where individuals are coordinators with possibly unique thresholds [granovetter1978threshold]. A decreasing payoff function, on the other hand, models the anticoordinators.
@@
Analyzing the outcome of these decision-making processes requires accounting for each individual’s actions, and consequently, a large body of research studied the steady state behavior of finite populations of best-responders. The problem has been investigated in various settings, ranging from well-mixed [zeroDeterminantS] to structured populations [ZHU2023110707, arefizadeh2023robustness, arditt2], and from time-invariant to time-varying thresholds [arditti2024robust]. Semi-tensor products were also utilized to analyze networked evolutionary games [guo2013algebraic, cheng2014finite] and logical dynamic games [boolean1, 10505780] .
@@
A finite population of heterogeneous coordinators, who differ in perceiving the commonality, admits equilibria [vanelli2020games] and equilibrates under asynchronous activation sequence [ramazi2020convergence]. A finite population of heterogeneous anticoordinators either equilibrates or converges to a two-state cycle [ramazi2017asynchronous, grabisch2020anti]. As for finite populations of coordinating and anticoordinating agents, it was shown that the dynamics may undergo perpetual fluctuations [grabisch2019model, grabisch2020anti, roohi]. The exact form and the conditions for the existence of the fluctuations appear to be an open problem.
@@
Researchers get around the challenges arising from the analysis of finite population dynamics by exploiting the associated deterministic mean dynamics [sandholm2010population, nowak2019homogeneous], where the population is assumed to be infinitely large. In this regard, most studies investigated the replicator dynamics which are the mean dynamics of the other popular update rule, imitation [comoImitation, replicator, replicator2, ramazireplicator]. Fewer studies considered the best-response mean dynamics, which are differential inclusions [hofbauer1995stability, golman2010basins, bestresponsePotential, berger2007two, theodorakopoulos2012selfish]. Some studies approximated the evolution of decisions in structured populations with ordinary differential equations [comoApproximation, ravazzi2023asynchronous, peidynamic].
@@
Although analysis of the mean dynamics is generally more straightforward compared to that of finite populations, the potential discrepancy between these two may question the validity of the approximation. Hence, much effort has been devoted to connecting the behavior of finite populations with the associated mean dynamics as the population size grows [benaim2005stochastic, benaim1998recursive, benaim2003deterministic, benaim2009mean]. The results on infinite horizon behavior connect the Birkhoff center of the mean dynamics with the support of the stationary measures of the Markov chain corresponding to the finite population’s dynamics [benaim1998recursive, roth2013stochastic].
@@
How can these results, which link the mean dynamics to finite population dynamics, be applied to the heterogeneous populations of coordinators and anticoordinators? To the best of our knowledge, no study has investigated the asymptotic behavior of infinite heterogeneous populations of all coordinators, all anticoordinators, or a mixture of both. Even if there were such results, could they shed light on the discrete population dynamics?
@@
We provide the answer in this paper, namely by using the deterministic differential inclusions as an approximation tool to reveal the asymptotic behavior of the discrete population dynamics as the population size approaches infinity. First, through an intuitive example, we show that the perpetual fluctuations do not scale as the population size grows. Second, we write the population dynamics as a Markov chain and obtain the associated continuous-time population dynamics–an upper semicontinuous differential inclusion. Third, we show that the collection of population dynamics Markov chains defines generalized stochastic approximation processes for the continuous-time population dynamics [roth2013stochastic]. This allows us to apply the results in [roth2013stochastic] to approximate the asymptotic behavior of the discrete population dynamics by finding the Birkhoff center of the continuous-time population dynamics, which we do in the last part of the paper.
@@
Our contribution is fourfold: i) We show that in the continuous-time population dynamics, the heterogeneous population of all coordinators may admit two types of equilibrium points, clean-cut and ruffled coordinator-driven, where clean-cut equilibria are asymptotically stable and ruffled are unstable–Lemma 2 and Theorem 2. On the contrary, the heterogeneous population of all anticoordinators admits only one globally asymptotically stable equilibrium point that could be either clean-cut or ruffled anti-coordinator driven–Lemma 2 and 1. These results are in line with the existing results on the steady-state behavior of finite heterogeneous populations of all coordinators [ramazi2020convergence] and all anticoordinators [ramazi2017asynchronous]. ii) We show that in the continuous-time population dynamics, the mixed heterogeneous population of coordinators and anticoordinators may admit all three types of equilibria, clean-cut, ruffled anticoordinator-driven, and ruffled coordinator-driven, where clean-cut and ruffled anticoordinator-driven equilibria are asymptotically stable, the other type is unstable, and the dynamics always converge to one of the equilibria–Lemma 2 and Theorem 2. iii) Consequently, we show that the reported perpetual fluctuations in the proportion of $\mathtt{A}$-players [roohi] almost surely vanish with the population size –Theorem 3 and 4. iv) Out of the analysis of the continuous-time population dynamics, insightful hints on the asymptotic behavior of the finite populations were obtained, such as having the same equilibria and similar long-term behavior. Hence, in general, for other population dynamics, it may prove useful to study the behavior of the continuous-time population dynamics, which is usually more straightforward, prior to the precise analysis of the finite population dynamics.
@@
The sets of real, nonnegative real, integer, nonnegative integer, and positive integer numbers are respectively shown by $\mathbb{R}$, $\mathbb{R}_{\geq 0}$, $\mathbb{Z}$, $\mathbb{Z}_{\geq 0}$, and $\mathbb{N}$. With an abuse of notation, for positive integers $n$ and $m$, define $\frac{1}{m}\mathbb{Z}^{n}=\{\frac{1}{m}\bm{y}\mid\bm{y}\in\mathbb{Z}^{n}\}$. For each $\in\mathbb{N}, [i][i$] denotes $\{1,2,\ldots,i\}$. Boldface letters refer to vectors. The notation $\bm{1}$ refers to a vector of appropriate size with all elements equal to $1$. The $i^{\text{th}}$ element of the vector $\mathbf{q}$ is indicated by ${q}_{i}$. The calligraphic font $\mathcal{X}$ represents a set. A sequence of variables $x_{0},x_{1},x_{2},\ldots$ is represented by $\langle x_{k}\rangle{{}_{k=0}^{\infty}}$. The Euclidean norm of a vector $\bm{x}$ is denoted by $|\bm{x}|$. By the notations $a_{1},b_{1}]-[a_{2},b_{2}]$ and $a_{1},b_{1}]-{c}$ we mean $a_{1}-a_{2},b_{1}-b_{2}]$ and $a_{1}-c,b_{1}-c]$, respectively. The notation $\text{cl}(\mathcal{X})$ denotes the closure of the set $\mathcal{X}$. A set-valued map ${\bm{\mathcal{V}}}(\bm{x})$ from $\mathbb{R}^{n}$ to $\mathbb{R}^{n}$ is denoted by notation ${\bm{\mathcal{V}}}:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$.
@@ II Problem Formulation
We consider a well-mixed population of $\mathsf{N}$ agents, labeled by $1,2,\ldots,\mathsf{N}$, choosing repeatedly between two strategies (or actions) $\mathtt{A}$ and $\mathtt{B}$ over a discrete time sequence $t\in\frac{1}{\mathsf{N}}\mathbb{Z}_{\geq 0}$. The time sequence is indexed by $k$ where $k=\mathsf{N}t$ (i.e., the index of time $\frac{k}{\mathsf{N}}$ is $k$). Each agent $i\in[\mathsf{N}]$ has a threshold $\tau(i)\in(0,1)$ and is either a coordinator (conformist) or an anticoordinator (a nonconformist). A coordinator $i$ (resp. an anticoordinator $i$) tends to choose strategy $\mathtt{A}$ whenever the population proportion of strategy-$\mathtt{A}$ players, excluding herself, denoted $x^{\mathsf{N}}_{-i}\in[0,1)$, does not fall short of (resp. does not exceed) her threshold, i.e., $x^{\mathsf{N}}_{-i}\geq\tau(i)$ (resp. $\leq\tau(i)$). The superscript $\mathsf{N}$ in $x^{\mathsf{N}}_{-i}$ denotes the population size.
@@
Coordinators (resp. anticoordinators) who have the same threshold build up a subpopulation, and there are altogether $\mathsf{p}^{\prime}$ (resp. $\mathsf{p}$) subpopulations of coordinators (resp. anticoordinators). Coordinating (resp. anticoordinating) subpopulations are labeled in the ascending (resp. descending) order of their thresholds by $1,\ldots,\mathsf{p}^{\prime}$ (resp. $\mathsf{p}$), that is, $\tau^{\prime}_{1}<\tau^{\prime}_{2}<\ldots<\tau^{\prime}_{\mathsf{p}^{\prime}}$ (resp. $\tau_{1}>\tau_{2}>\ldots>\tau_{\mathsf{p}}$), where $\tau^{\prime}_{p}$ (resp. $\tau_{p}$) is the threshold of the $p^{\text{th}}$ coordinating (resp. anticoordinating) subpopulation. So the heterogeneity of the population is captured by the distribution of population proportions over the total $\mathsf{p}+\mathsf{p}^{\prime}$ subpopulations as $\bm{\rho}=({\rho}_{1},\ldots,{\rho}_{\mathsf{p}},{\rho}^{\prime}_{\mathsf{p}^{\prime}},\ldots,{\rho}^{\prime}_{1})^{\top}$ where ${\rho}^{\prime}_{j}$ (resp. ${\rho}_{j}$) represents the ratio of the number of agents in coordinating (resp. anticoordinating) subpopulation $j$ to the population size $\mathsf{N}$. For convenience of vector indexing, for $j\geq\mathsf{p}+1$, we define ${\rho}_{j}={\rho}^{\prime}_{\mathsf{p}+\mathsf{p}^{\prime}+1-j}$. Note that there is no subpopulation of size zero, meaning that $\min_{j}\{\rho_{j}\}\geq\frac{1}{\mathsf{N}}$.
@@
According to the best-response update rule, the active agent chooses the strategy which maximizes her current payoff, i.e, $s_{i}(k+1)=\mathtt{A}$ if $u_{i}^{\mathtt{A}}(k)\geq u_{i}^{\mathtt{B}}(k)$ and $\mathtt{s}_{i}(k+1)=\mathtt{B}$, otherwise where $u_{i}^{\mathtt{j}}(k)$ is the payoff to agent $i$ from choosing strategy $\mathtt{j}\in\{\mathtt{A},\mathtt{B}\}$ at time index $k$. If the agents’ payoffs are affine functions of the population proportion of strategy-$\mathtt{A}$ players, update rules (1) and (2) become equivalent to the best-response update rule. See [ramazi2016networks] for more details.
@@
where $\mathit{x}^{\prime\mathsf{N}}_{p}$ (resp. $\mathit{x}^{\mathsf{N}}_{p}$) represents the proportion of $\mathtt{A}$-players who belong to coordinating (resp. anticoordinating) subpopulation $p$. We again use the notation $x^{\mathsf{N}}_{p}=x^{\prime\mathsf{N}}_{\mathsf{p}+\mathsf{p}^{\prime}+1-p}$ for $p>\mathsf{p}$. By defining the vector set
@@
The evolution of the state defines the population dynamics which are governed by update rules (1), (2), and the activation sequence of the agents. More specifically, the activation sequence is generated by the sequence of mutually independent random variables $\langle{A}_{k}\rangle{{}_{k=0}^{\infty}}$ where ${A}_{k}$ is the agent active at time index $k$ and takes values in $\mathsf{N}]$ with the distribution $\mathbb{P}[{A}_{k}=i]=\frac{1}{\mathsf{N}}$. A realization of an activation sequence $\langle{A}_{k}\rangle_{k=0}^{\infty}$ is persistent [ramazi2016networks], if for every agent $i$ and every $K>0$ there exists a time index $T>K$ such that $A_{T}=i$.
@@
The discrete best-response population dynamics are defined by the following discrete-time stochastic equation for $k=0,1,\ldots$
@@
where ${P}_{k}$ and ${S}_{k}$ are scalar random variables with distributions $\mathbb{P}[{P}_{k}=p]=\rho_{p}$, $\mathbb{P}[{S}_{k}=1|{P}_{k}=p]=x^{\mathsf{N}}_{p}/\rho_{p}$, $\mathbb{P}[{S}_{k}=2|{P}_{k}=p]=1-x^{\mathsf{N}}_{p}/\rho_{p}$, for $p\in[\mathsf{p}+\mathsf{p}^{\prime}]$, $\mathbf{e}_{{P}_{k}}$ is the ${P}_{k}$th standard basis vector in $\mathbb{R}^{\mathsf{p}+\mathsf{p}^{\prime}}$, and ${s^{*}}(p,\bm{x}^{\mathsf{N}},s)$ is a function where ${s^{*}}(p,\bm{x}^{\mathsf{N}},1)$ returns $1$ (resp. $2$) if $\mathtt{A}$ (resp. $\mathtt{B}$) is the preferred strategy of an $\mathtt{A}$-playing agent and ${s^{*}}(p,\bm{x}^{\mathsf{N}},2)$ returns $1$ (resp. $2$) if $\mathtt{A}$ (resp. $\mathtt{B}$) is the preferred strategy of a $\mathtt{B}$-playing agent in subpopulation $p$ at population state $\bm{x}^{\mathsf{N}}$ :
@@
[…]
@@
In Definition 1, the random variable ${P}_{k}$ is the subpopulation of the agent active at time index $k$, and the random variable ${S}_{k}$ equals $1$ (resp. $2$) if the strategy of the active agent at index $k$ equals $\mathtt{A}$ (resp. $\mathtt{B}$). Persistent realizations of the sequences of random variables $\langle P_{k}\rangle_{k=0}^{\infty}$ and $\langle S_{k}\rangle_{k=0}^{\infty}$ refer to their realizations under a persistent activation sequence. Function $s^{*}(p,\bm{x}^{\mathsf{N}},\mathtt{j})$ returns $1$ if $\mathtt{A}$ is the preferred strategy of an agent playing $\mathtt{j}$ in the subpopulation $p$ when the population is at state $\bm{x}^{\mathsf{N}}$, and it returns $2$ otherwise.
@@
In the context of public health, when a newly developed vaccine is introduced, some individuals choose to get vaccinated (i.e., strategy $\mathtt{A}$) as long as the vaccination coverage is below a certain threshold [ibuka2014free]. For them, this signals that not enough individuals are vaccinated, and they want to be immunized. Conversely, some individuals go for vaccination when a specific proportion of the population has already been vaccinated [SEANEHIA20172676] as then they perceive the vaccine as safe or a societal norm. The first group can be thought of as anticoordinators, while the latter are coordinators. The perception of what proportion of the population needs to be vaccinated to be considered rare or common varies among individuals, resulting in a heterogeneous population.
What is the asymptotic behavior of the population dynamics as the population size approaches infinity? Does the population state reach an equilibrium point where no agent tends to switch her current strategy? Or do the population dynamics undergo perpetual fluctuations?
We define a fluctuation set as a non-empty subset of the state space that is (i) positively invariant, meaning that once the state enters the set it never leaves it under any activation sequence, (ii) it is minimal, meaning that it does not admit a positively invariant proper subset, implying that once the state enters it, every state in the set will be visited under a persistent activation sequence, and (iii) it is non-singleton, implying that it is not an equilibrium. In what follows, we define it in terms of the dynamics in Definition 1.
@@
The dynamics of a finite mixed population consisting of both coordinating and anticoordinating subpopulations may undergo perpetual fluctuations [roohi], the characterization of which appears to be challenging and remains unsolved. Could the asymptotic behavior of discrete populations be revealed by investigating the associated mean dynamics, as the population size approaches infinity?